OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (−2, 1) and (4, 13)?

OpenStudy (anonymous):

@mohammad.arqum

OpenStudy (anonymous):

@mathavraj

OpenStudy (anonymous):

@skullpatrol @connor_

OpenStudy (anonymous):

try it yourself.. find the gradient... (y2-y1) / (x2-x1)

OpenStudy (anonymous):

i would if i could

OpenStudy (anonymous):

@connor_

OpenStudy (anonymous):

Find the slope first. m = (y2 - y1) / (x2 - x1) I'll do (13 - 1) / (4 + 2) 12 / 6 = 2 Either way when you subtract it, m = 2. Well now you can just substitute what you already have into the point slope form equation, which is y - y1 = m (x-x1) You can use either point, but you'll still get your answer as y - 1 = 2 ( x + 2)